OpenStudy (anonymous):

solve: 6x+4 3x+2 ----- = ------ 4x-3 2x+5 or, if this makes it easier to read (6x+4) / (4x-3) = (3x+2) / (2x+5)

OpenStudy (anonymous):

Cross multiply both sides, so you get \[(2x+5)(6x+4)=(3x+2)(4x-3)\] and then distribute \[12x^2+38x+20=12x^2-x-6\] move everything to one side \[39x+26=0\] and \[x=-26\div(39)\] Q.E.D